# Indirect Tunneling Enabled Spontaneous Time-Reversal Symmetry Breaking and Josephson Diode Effect in TiN/Al2O3/Hf0.8Zr0.2O2/Nb tunnel junctions

### v2

@@ I Introduction
The electrical control of the JD effect, which has so far been rarely achieved[27, 13], is also crucial for the applications of JD. Indeed, at the time of writing, only three studies reported magnetic-field-free JD effect in single JTJs (excluding non-tunneling or more complex devices), among which only one demonstrated the electrical control of the effect (Table 1). In addition, it is also desirable that JDs are fabricated using materials and processes compatible with the complementary metal-oxide semiconductor (CMOS) technology in view of its dominance in the current infrastructure anchoring semiconductor manufacturing. However, the fabrication of JDs by CMOS compatible techniques has not been demonstrated thus far. The use of CMOS-compatible atomic layer deposition (ALD) to prepare JTJs, which was attempted previously but not to achieve the JD effect[30, 31, 32], is highly desirable.
@@ III Quasiparticle tunneling
Figure 2: Quasiparticle tunneling and Josephson effect. (a) II-$V$ curves measured at various temperatures with critical and retrapping currents, $I_{\mathrm{c}}^{+,-}$ and $I_{\mathrm{r}}^{+,-}$, indicated. (b) dII/dVV vsvs. eVeV at various temperatures showing four peaks denoted by $\Delta_{\mathrm{1}}$, $\Delta_{\mathrm{2}}$, $\Delta_{\mathrm{3}}$, and $\Delta_{\mathrm{4}}$. These features are related to superconducting energy gaps of superconducting electrodes. (c) Values of the gap features with those obtained from the negative voltages marked by circles and those obtained from the positive voltages by triangles showing no nonreciprocal behavior. (d) $I_{\mathrm{c}}^{+,-}$ and $I_{\mathrm{r}}^{+,-}$ as a function of temperature (without the application of training current, see below). The error bars were obtained from the standard deviation (see text) with nonreciprocal behavior seen clearly.
The observation of two other peaks, $\Delta_{\mathrm{2}}$ and $\Delta_{\mathrm{4}}$, seen at voltages corresponding to the energy gaps of the two superconducting electrodes, $\Delta_{\mathrm{Nb}}$ and $\Delta_{\mathrm{TiN}}$, is not expected theoretically from direct quasiparticle tunneling between two superconductors[40]. However, such features were observed experimentally[43] and explained theoretically in a picture involving direct tunnelings of two quasiparticles at a finite bias voltage[44]. It should be noted that the Cooper pair tunneling facilitating the Josephson coupling corresponds to the coherent tunneling of two quasiparticles at the zero bias. The gap values were found to depend on the temperature as expected from the BCS theory (Fig. 2(c))[40]. Interestingly, because the gap value starts to drop appreciably when the temperature is raised to about 0.7Tc and the Tc values of Nb abd TiN are so different, the value of $\Delta_{\mathrm{3}}$ = $\Delta_{\mathrm{Nb}}$ - $\Delta_{\mathrm{TiN}}$ is seen to show an upturn from 2.1 to 3.2 K (Fig. 3(c)), as expected. The observations of the four gap features and the respective temperature dependences of them demonstrate that the tunnel barrier are pinhole-free even though the thickness varies spatially (which makes the double quasiparticle tunneling through thinnest and single one through less thin parts of the tunnel barrier competitive). This in turn demonstrates the power of ALD in the preparation of atomically thin tunnel barrier.
@@ IV Training of the JD effect
The averaged values of $I_{\mathrm{c}}^{+,-}$ and $I_{\mathrm{r}}^{+,-}$ shown in Fig. 2(d) reveal clearly nonreciprocal behavior (without the application of training current, see below). It is seen that the nonreciprocal behavior decreases as the temperature increases and disappears above around 3.2 K. A dip in $I_{\mathrm{c}}^{+}$ is also seen around 2.3 K. The temperature independence seen in $I_{\mathrm{r}}^{+,-}$, which is very different from that seen in $I_{\mathrm{c}}^{+,-}$, can be easily explained if the damping is dominated by high-frequency contributions rather than the frequency independent normal state junction resistance[45], which is likely to be the case in our experimental setup.
@@
In Fig. 3(a) and (c), it is shown that $I_{\mathrm{c}}^{-}$ is almost twice as large as $I_{\mathrm{c}}^{+}$ after applying a negative training current of -1.2 mA. After training with a positive current (+1.2 mA), the value of $I_{\mathrm{c}}^{-}$ was greatly reduced and $I_{\mathrm{c}}^{+}$ became almost twice as large as $I_{\mathrm{c}}^{-}$ (Fig. 3(b) and (d)). After the training, the diode efficiency was significantly enhanced. On the other hand, the effect of training disappears if the sample is warmed up above 10 K. Similar behavior in training was seen in additional samples featuring reduced diode efficiencies. Interestingly, the critical current flowing in the same direction as the training current appears to be unaffected or even slightly enhanced with its value close to the A-B limit. On the other hand, the critical current in the direction opposite to the training current is significantly suppressed. Furthermore, the training was found to fix the polarity of the JD as long as the JTJ is kept at low temperatures and below the larger critical current, which makes the device promising for nonvolatile memory in superconducting circuits.
@@ VI Spontaneous TRS breaking and electrical control of the JD effect
The simultaneous presence of positive and negative Josephson couplings in the JTJ will lead to the formation of supercurrent and corresponding magnetic flux loops (Fig. 4(c), which in turn leads to the TRS breaking, a necessary but not sufficient condition for the JD effect. To obtain the nonreciprocal $I_{\mathrm{c}}$ values, the distribution of the local Josephson coupling and its effect on the critical current of the junction need to be analyzed. Denoting the Josephson junction plane as the $x$-$y$ plane, we write the effective Hamiltonian in the form

### v3

@@ I Introduction
The electrical control of the JD effect, which has so far been rarely achieved[27, 13], is also crucial for the applications of JD. Only three studies reported magnetic-field-free JD effect in single JTJs (excluding non-tunneling or more complex devices), among which only one demonstrated the electrical control of the effect (Table 1). In addition, it is also desirable that JDs are fabricated using materials and processes compatible with the complementary metal-oxide semiconductor (CMOS) technology in view of its dominance in the current infrastructure anchoring semiconductor manufacturing. However, the fabrication of JDs by CMOS compatible techniques has not been demonstrated thus far. The use of CMOS-compatible atomic layer deposition (ALD) to prepare JTJs, which was attempted previously but not to achieve the JD effect[30, 31, 32], is highly desirable.
@@ III Quasiparticle tunneling
Figure 2: Quasiparticle tunneling and Josephson effect. (a) II-$V$ curves measured at various temperatures with critical and retrapping currents, $I_{\mathrm{c}}^{+,-}$ and $I_{\mathrm{r}}^{+,-}$, indicated. (b) dII/dVV vsvs. eVeV at various temperatures showing four peaks denoted by $\Delta_{\mathrm{1}}$, $\Delta_{\mathrm{2}}$, $\Delta_{\mathrm{3}}$, and $\Delta_{\mathrm{4}}$. These features are related to superconducting energy gaps of superconducting electrodes. (c) Values of the gap features with those obtained from the negative voltages marked by circles and those obtained from the positive voltages by triangles showing no nonreciprocal behavior. (d) $I_{\mathrm{c}}^{+,-}$ and $I_{\mathrm{r}}^{+,-}$ as a function of temperature without the application of training current. The error bars were obtained from the standard deviation (see text) with nonreciprocal behavior seen clearly.
[…]
@@ IV Training of the JD effect
The averaged values of $I_{\mathrm{c}}^{+,-}$ and $I_{\mathrm{r}}^{+,-}$ shown in Fig. 2(d) reveal clearly nonreciprocal behavior without the application of training current. It is seen that the nonreciprocal behavior decreases as the temperature increases and disappears above around 3.2 K. A dip in $I_{\mathrm{c}}^{+}$ is also seen around 2.3 K. The temperature independence seen in $I_{\mathrm{r}}^{+,-}$, which is very different from that seen in $I_{\mathrm{c}}^{+,-}$, can be easily explained if the damping is dominated by high-frequency contributions rather than the frequency independent normal state junction resistance[45], which is likely to be the case in our experimental setup.
@@
In Fig. 3(a) and (c), it is shown that $I_{\mathrm{c}}^{-}$ is almost twice as large as $I_{\mathrm{c}}^{+}$ after applying a negative training current of -1.2 mA. After training with a positive current (+1.2 mA), the value of $I_{\mathrm{c}}^{-}$ was greatly reduced and $I_{\mathrm{c}}^{+}$ became almost twice as large as $I_{\mathrm{c}}^{-}$ (Fig. 3(b) and (d)). After the training, the diode efficiency was significantly enhanced. On the other hand, the effect of training disappears if the sample is warmed up above 10 K. Similar behavior in training was seen in additional samples featuring reduced diode efficiencies. Interestingly, the critical current flowing in the same direction as the training current appears to be unaffected or even slightly enhanced with its value close to the A-B limit. On the other hand, the critical current in the direction opposite to the training current is significantly suppressed. The different diode effect onset temperatures for sample A and B, as shown in Fig. 3(c, d), may be due to differences in the spatial and energetic distribution of the localized states in the tunnel barrier, as will be discussed later. The retrapping current shows no diode effect, as shown in Fig. 3(c, d), most likely because the high-frequency damping originating from the leads dominates. Furthermore, the training was found to fix the polarity of the JD as long as the JTJ is kept at low temperatures and below the larger critical current, which makes the device promising for nonvolatile memory in superconducting circuits.
@@ VI Spontaneous TRS breaking and electrical control of the JD effect
The simultaneous presence of positive and negative Josephson couplings in the JTJ will lead to the formation of supercurrent and corresponding magnetic flux loops (Fig. 4(c), which in turn leads to the TRS breaking, a necessary but not sufficient condition for the JD effect. We note that our junctions are not $\phi_{0}$-junctions since a phase shift $\phi_{0}$ in the current phase relation would correspond to explicit rather than spontaneous TRS breaking [52, 53, 54, 7, 15]. To obtain the nonreciprocal $I_{\mathrm{c}}$ values, the distribution of the local Josephson coupling and its effect on the critical current of the junction need to be analyzed. Denoting the Josephson junction plane as the $x$-$y$ plane, we write the effective Hamiltonian in the form